OpenStudy (anonymous):

Please help me condense the following logarithm: (1/5)[log2 f + 3log2 g] - log2(3) - 6log2 (h-11) Thanks!!

OpenStudy (anonymous):

This item has common base of log2

OpenStudy (cwrw238):

use laws of logs log a + log b = log ab log a - log b = log a/b n log a = log a^n ill start you off: log2 f + 3 log2 g = log2 f + log2 g^3 = log2 fg^3

OpenStudy (anonymous):

i've gotten that far. I just don't understand how. I have the rules but can't seem to condensed or expand properly.

OpenStudy (cwrw238):

well next thing to do is change (1/5) log2 fg^3 - using the 3rd law of logs shown above - this is a sticky one i must admit, next apply 3rd law to the last term and then the 2nd law to last two terms

OpenStudy (anonymous):

\[z \le \log43+3\log6/2\log6??\]

OpenStudy (anonymous):

is this correct?

OpenStudy (mertsj):

Do you need help with this problem?